OpenStudy (anonymous):

expand logarithm log b 4 x 2 y

OpenStudy (solomonzelman):

\[\log~(b^4x^2y)\]

OpenStudy (solomonzelman):

\[Log(xy)=Log~x~~+~~Log~y\]\[Log~(a^b)~=~b~Log~a\]

OpenStudy (solomonzelman):

use these rules (in order) to expand this logarithm.

OpenStudy (solomonzelman):

Use this example. \[\color{blue} {Log~(s^3~d^2)}\]\[\color{blue} {Log~(s^3)}\color{blue} {~+~Log~(d^2)}\]\[\color{blue} {3~Log~(s)}\color{blue} {~+~2~Log~(d)}\]

OpenStudy (solomonzelman):

\[\color{purple} {\log b^4x^2 y}\]\[\color{purple} {\log (b^4)}+\color{purple} {\log (x^2)}+\color{purple} {\log (y)}\]\[\color{purple} {4~\log (b)}+\color{purple} {\log (2~x)}+\color{purple} {\log (y)}\]here, but in the future please refer to the helper's tips.